In a classroom, there are 2 more columns than rows of desks. If there are 48 students find how the desks are arrange.
Answer by cleomenius(959)

(x) (x +2) = 48
x^2 + 2x - 48 = 0
(x -6)(x + 8)
x = 6, x = -8
using 6, x = 6 and x + 2 = 8.
6 * 8 = 48.
Cleomenius.